# Multimessenger Astronomy Beyond the Standard Model: New Window from Quantum Sensors

### v1

@@ Appendix B Sensitivity for Additional Couplings
Figure 22: Sensitivity to ALP linear coupling to photons, $g_{\phi\gamma\gamma}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-20}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-15}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents parameter space excluded by Chandra observations of spectral distortions from $\gamma$-$\phi$ oscillations in luminous quasar H1821+643 sisk2022new . The black line shows the projected sensitivity of future experiments with $g_{\rm DM}=3\times 10^{-16}~{}\mathrm{GeV}^{-1}$ and $t_{\rm int,DM}=1~{}\mathrm{year}$, based on proposed superconducting cavities berlin2021heterodyne and lumped-element detectors dmradio2022projected , scaled using Eq. (VI.2). The black hatched and brown regions are excluded by ALP energy constraints and the uncertainty principle. Dashed purple and red lines indicate detection delays of 1 day and 1 year relative to a light-speed signal. For these benchmarks, no viable parameter space exists for multimessenger astronomy with ALPs.

### v2

@@ Appendix B Sensitivity for Additional Couplings
Figure 22: Sensitivity to ALP linear coupling to photons, $g_{\phi\gamma\gamma}$, as a function of the ULB scalar energy $\omega$, considering a relativistic ULB field burst from a source located at $R=10~{}\mathrm{kpc}$ from Earth, releasing a total energy of $E_{\rm total}=10^{-2}M_{\odot}$ in the form of ULBs. Benchmark ULB masses of $m_{\phi}=10^{-20}~{}\mathrm{eV}$ (left column) and $m_{\phi}=10^{-15}~{}\mathrm{eV}$ (right column), with intrinsic burst times of $t_{*}=1~{}\mathrm{s}$ (top row) and $t_{*}=100~{}\mathrm{s}$ (bottom row) are displayed. The gray region represents parameter space excluded by Chandra observations of spectral distortions from $\gamma$-$\phi$ oscillations in luminous quasar H1821+643 sisk2022new . The black line shows the projected sensitivity of future experiments with $g_{\rm DM}=3\times 10^{-16}~{}\mathrm{GeV}^{-1}$ and $t_{\rm int,DM}=1~{}\mathrm{year}$, based on proposed superconducting cavities berlin2021heterodyne and lumped-element detectors dmradio2022projected , scaled using Eq. (VI.2). The black hatched and brown regions are excluded by ALP energy constraints and the uncertainty principle. Dashed purple and red lines indicate detection delays of 1 day and 1 year relative to a light-speed signal. The green region denotes viable parameter space for multimessenger astronomy with scalars, considering all constraints and benchmarks.